Find the first six terms of the sequence.
a1 = -8, an = 5 • an-1 (1 point)
choices are below
-8, -40, -200, -1000, -5000, -25,000

-8, -40, -35, -30, -25, -20

0, 5, -40, -35, -30, -25

-40, -200, -1000, -5000, -25,000, -125

Answer by mathsolverplus(88)

The correct choice would be the first one.
a1 = -8
a2 = 5*a1 = 5*-8 = -40
a3 = 5*a2= 5*-40 = -200


If this answer has helped you and you need help on additional problems, please follow us on Instagram or message us via social media! We also help with SAT Math Questions!

***Instagram (@mathsolver.plus): https://www.instagram.com/mathsolver.plus/
Twitter (@MathSolverPlus): https://twitter.com/MathSolverPlus
Snapchat (@mathsolverplus)
Email: mathsolverplus@gmail.com